A physics geek walks 16m(n) in 10s and the 10m in 5s calculate the a) distance b) displacement c) speed d) velocity
1
Expert's answer
2018-09-17T15:21:08-0400
a)
d=16+10=26 m.
b)
d=16+10=26 m due north.
c)
v=26/(10+5)=1.7 m/s.
d)
v=26/(10+5)=1.7 m/s due north.